Find an equation of the line having the given slope and containing the given point.
m=-9,(2,4)

Answer by ankor@dixie-net.com(22740)

Find an equation of the line having the given slope and containing the given point.
m=-9,(2,4)
:
Using the point/slope formula: y - y1 = m(x - x1)
y - 4 = -9(x - 2)
y - 4 = -9x + 18
y = -9x + 18 + 4
y = -9x + 22